4. In a right triangle find the length of the side not given
b=2 and c=√13

Answer by nyc_function(2741)

Use the Pythagorean Theorem.
a^2 + b^2 = c^2
a^2 + (2)^2 = (√13)^2
We are searching for side a.
a^2 + 4 = 13
a^2 = 13 - 4
a^2 = 9
Take the square root of both sides.
sqrt{a^2} = sqrt{9}
a = 3
Done!